Solve the following using elimination:
y=-2x +5
3x-2y=4

Answer by Alan3354(69443)

Solve the following using elimination:
y=-2x +5
3x-2y=4
------------
rearrange into standard form
2x + y = 5 eqn 1
3x -2y = 4 eqn 2
------------
Multiply eqn 1 by 2
4x +2y = 10 eqn 1
3x -2y = 4 eqn 2
Add the 2 eqns
7x +0y = 14
x = 2
-------
Sub into eqn 1
2*2 + y = 5
4 + y = 5
y = 1
--------
check, they both work